OpenStudy (anonymous):

(PLEASE TELL ME HOW TO SOLVE IT DON'T JUST DO IT FOR ME...) How many sides does an unnamed polygon have if the sum of the measures of the interior angles is 2340 degrees?

OpenStudy (anonymous):

No one is just going to do your homework for you someone is going to give an explanation.

OpenStudy (anonymous):

(n-2)180=sum of the interior angles. Where n=number of angles of a regular polygon. We know that the sum of the interior angles is 2340 so (n-2)180=2340 divide both sides by 180 and get n-2=13 add 2 to both sides and get n=15 Therefore, this polygon has 15 sides.

OpenStudy (anonymous):

I know your just copied that off a website.....that doesn't help me because I already looks at that site and I don't know where you're getting these numbers so.....

OpenStudy (anonymous):

HaHaHa great deduction. Well you have the answer you wanted right? lol I know I couldn't answer it with my own brain I'm only in 9th grade.

OpenStudy (anonymous):

And I'm in 7th your point ...?

OpenStudy (anonymous):

lol I don't remember learning that....

OpenStudy (anonymous):

Well whatever cause I just wanted to know how to solve and I still don't get it so whatever

OpenStudy (anonymous):

My point is Algebra 1 is so hard you forget stuff like this.

OpenStudy (anonymous):

Well i'am in 7th and I'm no learning this.

OpenStudy (anonymous):

I'm in Pre Algebra

OpenStudy (anonymous):

(not) i mean

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Nice pre-algebra is sooo not hard at all.

OpenStudy (anonymous):

I know but many just hate it in my class btw this is how my class looks

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

7th grade isn't that hard either.

OpenStudy (anonymous):

Then why can't you answer it...if you don't remember why bother try answering like rlly

OpenStudy (anonymous):

I can't answer it because I forgot that's why.

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

You forget there is like a whole 2 years now since I've been in 7th grade.

OpenStudy (anonymous):

If it's so easy why don't you remember the easy things

OpenStudy (anonymous):

yep XD he doesn't like that.

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

lol CAUSE IT'S BEEN 2 YEARS!!!!

OpenStudy (jdoe0001):

\(\bf \textit{sum of interior angles of a polygon}=180({\color{red}{ n}}-2)\qquad {\color{red}{ n}}=\textit{number of sides} \\ \quad \\ thus\qquad 2340=180({\color{red}{ n}}-2)\) solve for "n", to get the number of sides in the polygon

OpenStudy (anonymous):

I don't get where you got the # 180 and the # 2

OpenStudy (anonymous):

It's just the formula.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thank you jdoe0001

OpenStudy (anonymous):

The formula for the sum of the interior angles of a shape with n sides is: 180 * (n-2) = S Substitute in 2340 for S and solve for n: 180 * (n-2) = 2340 Divide both sides by 180: (n-2) = 13 Add 2 to both sides: n = 15 The polygon has 15 sides

OpenStudy (anonymous):

hope this helped :)

OpenStudy (anonymous):

thanks now i understand thanks mjoxxprm360

OpenStudy (anonymous):

anytime

OpenStudy (anonymous):

also guys one more thing

OpenStudy (anonymous):

http://www.hurr-durr.com/

OpenStudy (anonymous):

XD